OpenStudy (anonymous):

Next question, Calculate the enthalpy change per mole of oxygen consumed in the reaction: 2U(s) + 3O2(g) → 2UO3(s). I know I'm supposed to use the CRC Handbook to find enthalpy values, but I'm not quite sure what to do after that...

OpenStudy (anonymous):

Since this reaction is the standard heat of formation for UO3(s) you need to find the value for the enthalpy of formation from your table. I found a value of -1223.8 in a google search (use the value from your authorized source). Now , this represents the amount of heat released when 1 mol of UO3(s) is formed. To form 1 mol of UO3(s) we need 3/2 moles of O2(g). Your question asks for the heat of formation of UO3(s) per mole of O2(g) consumed. This would be:\[-1223.8 kJ/mol UO3 * \frac{2 mol O2}{3 mol UO3}=-815.7 \frac{kJ}{mol O_{2}(g)}\]at standard temperture and pressure conditions.

OpenStudy (anonymous):

One way to see this is to re-write the reaction as:\[U(s)+\frac{3}{2}O_{2}(g)\rightarrow UO_{3}(s)\]My value of -1223.8 kJ is for this reaction. Now, to get the amount of heat released per mol of O2, multiply the equation through by 2/3 including the value -1223.8 kJ.

OpenStudy (anonymous):

Thank you! It worked!